Water is flowing in a hose with a velocity of 1.5 m/s and a pressure of 3 kPa. At

    the nozzle the pressure decreases to atmospheric pressure (101300 Pa), there

    is no change in height. Use the Bernoulli equation to calculate the velocity of the

   water exiting the nozzle. The density of water is 1000 kg/m3 and gravity g is 9.8 

    m/s2.


1
Expert's answer
2022-06-15T07:03:50-0400

"\frac{\rho }2(v_2^2-v_1^2)=p_2-p_1,"

"v_1=\sqrt{v_2^2-\frac{2(p_2-p_1)}{\rho}}=0.8~\frac ms."